# Deblending the MIGHTEE-COSMOS survey with XID+: The resolved radio source counts to Jy

### v1

@@ 2.3 Star-Galaxy Classification
A significant challenge in constructing a prior list from deep optical/IR catalogs is the presence of stellar sources. Since most stars are not detectable radio sources (e.g Driessen2024; Okwudili2025), their inclusion in the prior list would introduce a population of non-emitting sources. This could cause the deblending algorithm to assign radio flux incorrectly from nearby galaxies to these stellar positions, thereby contaminating the measurements of genuine radio galaxies. Therefore, a reliable method for star-galaxy separation is an important preliminary step.
@@ 2.3.1 Methodology
Our star-galaxy classification process begins with the 650,148 sources matched between our master catalog and COSMOS2020. We use the HSC PDR3 photometry for classification because its improved depth enables more reliable morphological separation compared to earlier releases. We apply two selection criteria to ensure that we use robust photometry. First, we select sources with a signal-to-noise ratio in the Ks-band of SNR > 5, reducing the sample to 506,136 sources.
@@
To classify these sources, we use a two-stage unsupervised machine learning process, following Cook2024. Unsupervised learning has the advantage of not being biased by a given training set, which is critical for faint source populations where complete and representative training data are unavailable (see Fotopoulou2024, for a review). First, we construct a high-dimensional feature space from the photometric data. We use magnitudes from the optical HSC survey and the near-infrared from VISTA to generate 28 color combinations. These colors, combined with the 8 original magnitude measurements (), serve as 36 features for each source. While some color features are mathematically redundant, we include all of them, as dimensionality reduction techniques like Uniform Manifold Approximation and Projection (UMAP; McInnes2020) effectively extracts the most important underlying information from complete feature sets (Cook2024). Furthermore, we verified that a reduced set of independent features (e.g., a single magnitude band plus n-1 independent colors) produced a consistent classification, confirming the robustness of the approach. This high-dimensional feature space provides a more robust basis for distinguishing between stellar and galactic populations than is possible in any single two-dimensional parameter space (e.g Strateva2001; Hewett2006).
@@
Before applying the machine learning algorithms, the feature set is first cleaned by removing sources with incomplete photometry, which yields a working sample of 492,741 sources. This feature set is then scaled using the StandardScaler package from scikit-learn (Pedregosa2011). This process applies a standard Z-score transformation, rescaling each feature to have a mean of zero and a standard deviation of one, which ensures that all features contribute equally to the analysis.
@@
The first stage of the process applies the UMAP algorithm to this feature space (McInnes2020). UMAP is a non-linear dimensionality reduction technique that reduces the 36-dimensional feature space to a 10-dimensional embedding111These 10 output dimensions are abstract representations; each is a complex, non-linear combination of the original 36 photometric features and does not correspond directly to any single physical property like a specific color or magnitude. while preserving the essential topological structure of the data (Becht2019). Its ability to capture the complex, non-linear relationships inherent in multi-band photometric data makes it a superior choice to simpler linear methods like Principal Component Analysis (PCA; Jolliffe2016). Furthermore, UMAP has been shown to be more scalable and better at preserving the global structure of the data, crucial for separating the main star and galaxy clusters, than comparable techniques (McInnes2020).
@@
The second stage applies the Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) algorithm (McInnes2017) to identify distinct clusters in the UMAP-reduced space. HDBSCAN is a density-based algorithm that can find clusters of varying shapes and densities without requiring the number of clusters to be specified beforehand, making it effective at isolating the primary stellar and galactic populations (e.g. Logan2020; Fotopoulou2024).
@@ 2.3.2 Application
We applied this methodology to the filtered sample of sources derived from the cross-matched dataset. The process successfully separates the sources, yielding a catalog of high-confidence galaxies ( per cent) and stars (4.7 per cent), as detailed in Table 1. The resulting UMAP projection, which clearly shows the distinct clusters, is presented in Figure 1. The classification is visualized on the traditional vs. color-color diagram (e.g., Jarvis2013) in Figure 2. In this two-dimensional projection, while the bulk of the populations lie on distinct loci, the contours reveal a region of degeneracy where stars and galaxies overlap. Traditional methods, which rely on defining a fixed boundary or simple curves relative to the stellar locus (e.g., Jarvis2013), inevitably suffer from contamination or incompleteness in this overlapping regime. However, our machine learning approach uses the full high-dimensional feature space to resolve these degeneracies and successfully disentangles stars and galaxies that are indistinguishable in the vs. plane alone.
@@ 2.4 The Galaxy Sample and prior list construction
Following star-galaxy classification, we define a galaxy sample from which deblending priors are constructed. We select all galaxies from our filtered sample that have a total infrared luminosity () estimate from the COSMOS2020 catalog. This results in a galaxy list of 331,442 galaxies within the region of overlapping deep multi-wavelength coverage. This sample forms the basis for all subsequent prior list construction.
Constructing a prior list presents a trade-off between completeness and purity. An inclusive prior ensures that all potential radio-emitting galaxies are considered, but at the risk of including a large number of undetectable radio sources. As demonstrated by Hurley2017, an excess of such priors can degrade the flux measurements of genuine detections by incorrectly partitioning the map’s flux density. To investigate this trade-off and determine the optimal strategy for MIGHTEE, we construct two distinct prior lists from our galaxy sample, each designed to test a different deblending philosophy.
The first prior list is designed to be as inclusive as possible. It incorporates all 331,442 sources from our galaxy sample that have an measurement. We supplement this list with 4,754 sources detected above 3 in the VLA-COSMOS 3 GHz survey (Smolcic2017), as extracted by Jin2018, specifically adding those that do not have counterparts in our optical/NIR sample. This results in a candidate list of 336,196 unique sources. This Comprehensive LIR Prior represents a compromise, aiming for high completeness by including any source with evidence of significant star formation or AGN activity (via ) without being as exhaustive as including all 470,000 galaxies in the master sample.
@@
First, we predict the 1.4 GHz radio flux density for each of the 331,442 galaxies in our galaxy sample based on their estimated infrared luminosity, using the well-established IRRC (Helou1985; Condon1992; Yun2001). The methodology is as follows:
@@
The rest-frame 1.4 GHz radio luminosity () is calculated from the total infrared luminosity () using the linear correlation from CookRHW2024, given by .
@@
This radio luminosity is converted to an expected observed flux density () using the source’s redshift () and corresponding luminosity distance (), including a k-correction of the form . We assume a spectral index , a typical value for the synchrotron-dominated emission of star-forming galaxies at these frequencies (Condon1992).
@@
The results of this prediction step are shown in Figure 3. As expected, the 98,825 sources predicted to be ‘Definitely Detectable’ occupy the high- regime across all redshifts, with the 2,353 ‘Possibly Detectable’ sources forming a boundary region where detection is plausible given uncertainties.
@@ 3 XID+ probabilistic deblending technique
Having established a prior list of source positions from ancillary data, we now address the challenge of measuring their flux densities in the confused MIGHTEE radio maps. We use the probabilistic deblending framework XID+ (Hurley2017) to perform this task. Even among prior-driven deblending approaches, many rely on maximum-likelihood estimation to assign flux (e.g., FASTPHOT, Bethermin2010). While computationally efficient, these approaches can struggle in highly confused regimes, often incorrectly assigning the total flux density of a blend to a single source and failing to provide robust uncertainty estimates (e.g. Roseboom2010; Wang2014).
XID+ overcomes these limitations by adopting a fully probabilistic Bayesian framework (Hurley2017). Instead of seeking a single best-fit solution, XID+ explores the entire parameter space to map out the full posterior probability distribution for the flux of every source (Hurley2017). This approach has several advantages. First, it provides uncertainty quantification by naturally capturing the credible intervals and degeneracies that arise when multiple sources are close together. Second, the joint posterior reveals covariance between sources, quantifying how the flux density estimate of one source is correlated with its neighbors. Third, the Bayesian framework correctly models the behavior of sources at or below the noise level.
The core assumption of XID+ is that the observed map data, , can be modelled as a linear combination of the flux densities, , of a known set of prior sources, each treated as a point source convolved with the instrument’s Point Spread Function (PSF), plus a background () and noise component (). For this work, we use the Gaussian restoring beam from the low-resolution MIGHTEE data. This noise is modelled as a Gaussian that accounts for both thermal noise and confusion noise arising from the blending of faint unresolved sources (e.g. Condon1974). For a map flattened into a vector of pixels, this can be expressed as:
@@
Here, is the likelihood of observing the data given a set of fluxes, and is the prior probability distribution of the fluxes. The likelihood is modelled as a multivariate Gaussian distribution based on the noise properties of the map. The standard XID+ framework calculates the total variance by adding the thermal variance (from a user-provided map) and a fitted confusion variance in quadrature (). However, a pure thermal noise map was not available for the MIGHTEE DR1 data products; instead, a total noise RMS map, which already accounts for both effects, was provided (Hale2025). This means the standard likelihood calculation within XID+ could not be used without modification.
@@
For the deblending performed in this work, we use a simple, uninformative (flat) prior for the flux of each source, typically bounded between 0 and a sufficiently high value, to avoid biasing the resulting measurements towards any pre-conceived model (Hurley2017; Wang2024). The role of the ancillary data is therefore to define an optimal and pure list of source positions, rather than to inform the flux priors directly.
@@
Solving for the full posterior distribution in Equation 2 is a high-dimensional problem that is computationally intractable to solve analytically. XID+ addresses this by using the probabilistic programming language Stan (Carpenter2017). Specifically, it uses the No-U-Turn Sampler (NUTS, Hoffman2014), to efficiently draw samples from the high-dimensional posterior distribution. To accommodate the MIGHTEE DR1 noise map, we modified the XID+ Stan code to use the pre-computed total variance from the DR1 RMS map at each pixel () directly in the likelihood calculation. While this means we do not derive a separate confusion noise component, it ensures the deblending is performed using the most accurate, position-dependent noise estimate available. An important parameter governing the sampler’s exploration is max_treedepth, which controls the complexity of the integration paths, with a default value of 10. Testing on real MIGHTEE data in Section 5.3 shows that increasing this to 15 was required for convergence. Following the recommendation of Hurley2017, we run four independent MCMC chains to ensure convergence.
@@
Fitting the whole map at once is computationally unfeasible, so the analysis is partitioned into small tiles using the HEALPix framework (Gorski2005), each 3.4 arcmin across (HEALPix order 10). However, to ensure an accurate flux measurement for sources within one of these small tiles, the deblending fit must account for flux spilling in from bright sources in the surrounding area. Therefore, the fit is performed on a much larger ‘fitting’ tile, 27.5 arcmin across (HEALPix order 7), which fully encompasses the smaller results tile. Once the fit on the larger tile is complete, only the results for sources within the central small tile are retained. The primary output for each source within a successfully fitted tile is a chain of samples from its posterior distribution. The point estimates for the deblended flux densities for each source are taken as the median (50th percentile) of its posterior distribution for each source. The final catalog is produced by repeating this process for all tiles across the map and combining the results.

### v2

@@ 2.3 Star-Galaxy Classification
A significant challenge in constructing a prior list from deep optical/IR catalogues is the presence of stellar sources. Since most stars are not detectable radio sources (e.g Driessen et al. 2024; Egbo et al. 2025), their inclusion in the prior list would introduce a population of non-emitting sources. This could cause the deblending algorithm to assign radio flux incorrectly from nearby galaxies to these stellar positions, thereby contaminating the measurements of genuine radio galaxies. Therefore, a reliable method for star-galaxy separation is an important preliminary step.
@@ 2.3.1 Methodology
Our star-galaxy classification process begins with the 650,148 sources matched between our master catalogue and COSMOS2020. We use the HSC PDR3 photometry for classification because its improved depth enables more reliable morphological separation compared to earlier releases. We apply two selection criteria to ensure that we use robust photometry. First, we select sources with a signal-to-noise ratio in the Ks-band of SNR > 5, reducing the sample to 506,136 sources.
@@
To classify these sources, we use a two-stage unsupervised machine learning process, following Cook et al. (2024b). Unsupervised learning has the advantage of not being biased by a given training set, which is critical for faint source populations where complete and representative training data are unavailable (see Fotopoulou, 2024, for a review). First, we construct a high-dimensional feature space from the photometric data. We use magnitudes from the optical HSC survey and the near-infrared from VISTA to generate 28 colour combinations. These colours, combined with the 8 original magnitude measurements (), serve as 36 features for each source. While some colour features are mathematically redundant, we include all of them, as dimensionality reduction techniques like Uniform Manifold Approximation and Projection (UMAP; McInnes et al., 2018) effectively extracts the most important underlying information from complete feature sets (Cook et al., 2024b). Furthermore, we verified that a reduced set of independent features (e.g., a single magnitude band plus n-1 independent colours) produced a consistent classification, confirming the robustness of the approach. This high-dimensional feature space provides a more robust basis for distinguishing between stellar and galactic populations than is possible in any single two-dimensional parameter space (e.g Strateva et al. 2001; Hewett et al. 2006).
@@
Before applying the machine learning algorithms, the feature set is first cleaned by removing sources with incomplete photometry, which yields a working sample of 492,741 sources. This feature set is then scaled using the StandardScaler package from scikit-learn (Pedregosa et al., 2011). This process applies a standard Z-score transformation, rescaling each feature to have a mean of zero and a standard deviation of one, which ensures that all features contribute equally to the analysis.
@@
The first stage of the process applies the UMAP algorithm to this feature space (McInnes et al., 2018). UMAP is a non-linear dimensionality reduction technique that reduces the 36-dimensional feature space to a 10-dimensional embedding111These 10 output dimensions are abstract representations; each is a complex, non-linear combination of the original 36 photometric features and does not correspond directly to any single physical property like a specific colour or magnitude. while preserving the essential topological structure of the data (Becht et al., 2019). Its ability to capture the complex, non-linear relationships inherent in multi-band photometric data makes it a superior choice to simpler linear methods like Principal Component Analysis (PCA; Jolliffe and Cadima 2016). Furthermore, UMAP has been shown to be more scalable and better at preserving the global structure of the data, crucial for separating the main star and galaxy clusters, than comparable techniques (McInnes et al., 2018).
@@
The second stage applies the Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) algorithm (McInnes et al., 2017) to identify distinct clusters in the UMAP-reduced space. HDBSCAN is a density-based algorithm that can find clusters of varying shapes and densities without requiring the number of clusters to be specified beforehand, making it effective at isolating the primary stellar and galactic populations (e.g.,Logan and Fotopoulou 2020; Fotopoulou 2024).
@@ 2.3.2 Application
We applied this methodology to the filtered sample of sources derived from the cross-matched dataset. The process successfully separates the sources, yielding a catalogue of high-confidence galaxies ( per cent) and stars (4.7 per cent), as detailed in Table 1. The resulting UMAP projection, which clearly shows the distinct clusters, is presented in Figure 1. The classification is visualized on the traditional vs. colour-colour diagram (e.g., Jarvis et al., 2013) in Figure 2. In this two-dimensional projection, while the bulk of the populations lie on distinct loci, the contours reveal a region of degeneracy where stars and galaxies overlap. Traditional methods, which rely on defining a fixed boundary or simple curves relative to the stellar locus (e.g., Jarvis et al., 2013), inevitably suffer from contamination or incompleteness in this overlapping regime. However, our machine learning approach uses the full high-dimensional feature space to resolve these degeneracies and successfully disentangles stars and galaxies that are indistinguishable in the vs. plane alone.
@@ 2.4 The Galaxy Sample and Prior List Construction
Following star-galaxy classification, we define a galaxy sample from which deblending priors are constructed. We select all galaxies from our filtered sample that have a total infrared luminosity () estimate from the COSMOS2020 catalogue. This results in a galaxy list of 331,442 galaxies within the region of overlapping deep multi-wavelength coverage. This sample forms the basis for all subsequent prior list construction.
Constructing a prior list presents a trade-off between completeness and purity. An inclusive prior ensures that all potential radio-emitting galaxies are considered, but at the risk of including a large number of undetectable radio sources. As demonstrated by Hurley et al. (2017), an excess of such priors can degrade the flux measurements of genuine detections by incorrectly partitioning the map’s flux density. To investigate this trade-off and determine the optimal strategy for MIGHTEE, we construct two distinct prior lists from our galaxy sample, each designed to test a different deblending philosophy.
2.4.1 Comprehensive Prior
The first prior list is designed to be as inclusive as possible. It incorporates all 331,442 sources from our galaxy sample that have an measurement. We supplement this list with 4,754 sources detected above 3 in the VLA-COSMOS 3 GHz survey (Smolčić et al., 2017), as extracted by Jin et al. (2018), specifically adding those that do not have counterparts in our optical/NIR sample. This results in a candidate list of 336,196 unique sources. This Comprehensive LIR Prior represents a compromise, aiming for high completeness by including any source with evidence of significant star formation or AGN activity (via ) without being as exhaustive as including all 470,000 galaxies in the master sample.
2.4.2 The Radio-Likely Prior
@@
First, we predict the 1.4 GHz radio flux density for each of the 331,442 galaxies in our galaxy sample based on their estimated infrared luminosity, using the well-established IRRC (Helou et al., 1985; Condon, 1992; Yun et al., 2001). The methodology is as follows:
@@
The rest-frame 1.4 GHz radio luminosity () is calculated from the total infrared luminosity () using the linear correlation from Cook et al. (2024a), given by .
@@
This radio luminosity is converted to an expected observed flux density () using the source’s redshift () and corresponding luminosity distance (), including a k-correction of the form . We assume a spectral index , a typical value for the synchrotron-dominated emission of star-forming galaxies at these frequencies (Condon, 1992).
@@
The results of this prediction step are shown in Figure 3. As expected, the 98,825 sources predicted to be ‘Definitely Detectable’ occupy the high- regime across all redshifts, with the 2,353 ‘Possibly Detectable’ sources forming a boundary region where detection is plausible given uncertainties. We note that the tail of ‘Definitely Detectable’ sources extending to very high redshift in Figure 3 should be interpreted with caution due to the lack of FIR constraints in the SED fitting. Furthermore, physical effects, such as increased inverse-Compton cooling off the Cosmic Microwave Background (CMB) photons, are expected to suppress the observed synchrotron emission at these redshifts (e.g., Murphy, 2009; Lacki and Thompson, 2010; Wu et al., 2017; Whittam et al., 2025). However, these potentially spurious high-redshift sources constitute a negligible fraction () of the full prior list and therefore do not impact the derived source counts or deblending results.
@@
2.4.3 Validation of Prior Selection
To validate the reliability of our SED-derived selection, particularly given the potential uncertainties in the energy-balance estimates which lack direct FIR constraints, we compare the IRRC-selected subset of our prior list to the updated Super-deblended catalogue (Jin et al., 2018; Sillassen et al., 2024). This catalogue serves as a valuable independent reference because its prior list was constructed using a different approach. While our method selects sources based on their predicted radio flux derived from , the Super-deblended prior list (containing 200,000 sources) is a stellar mass-limited sample. It was constructed by selecting sources from the COSMOS2020 catalogue (Weaver et al., 2022), the VLA 3 GHz survey (Smolčić et al., 2017), and the A3COSMOS catalogue (Liu et al., 2019). Since stellar mass correlates strongly with star formation rate (e.g., Noeske et al., 2007; Whitaker et al., 2012), and radio luminosity is a direct tracer of star formation, our SED-derived prior and the Jin et al. (2018) stellar-mass limited approach should target essentially the same underlying galaxy population. Any significant differences would therefore indicate potential issues with our prior list fidelity.
We cross-match our prior list with the Super-deblended catalogue using a search radius of 0.3 arcsec. We restrict this analysis to the clean area defined by the FLAG_COMBINED flag and the stellar mask (detailed in Section 5). This filtering reduces our galaxy sample to 298,894 sources. Within this area, our IRRC-based method identifies a Radio-Likely Prior subset of 87,259 sources. When matched against the super-deblended catalogue, we find that per cent of our sources have counterparts in the independent catalogue. This near-perfect agreement with a high-fidelity benchmark confirms that our -based selection effectively isolates the detectable radio population that follows the IRRC. Furthermore, it demonstrates that the uncertainties in arising from the lack of FIR data in the SED fitting do not significantly compromise the selection of the radio-emitting sample.
Having validated the high purity of the Radio-Likely Prior against independent data, we now turn to simulations to rigorously test the deblending accuracy. The relative performance of these two prior lists, one designed for inclusivity and the other for purity, will be rigorously tested in Section 4 to determine the optimal strategy for our final scientific analysis.
@@ 3 XID+ probabilistic deblending technique
Having established a prior list of source positions from ancillary data, we now address the challenge of measuring their flux densities in the confused MIGHTEE radio maps. We use the probabilistic deblending framework XID+ (Hurley et al., 2017) to perform this task. Even among prior-driven deblending approaches, many rely on maximum-likelihood estimation to assign flux (e.g., FASTPHOT, Béthermin et al. 2010). While computationally efficient, these approaches can struggle in highly confused regimes, often incorrectly assigning the total flux density of a blend to a single source and failing to provide robust uncertainty estimates (e.g.,Roseboom et al. 2010; Wang et al. 2014).
XID+ overcomes these limitations by adopting a fully probabilistic Bayesian framework (Hurley et al. 2017). Instead of seeking a single best-fit solution, XID+ explores the entire parameter space to map out the full posterior probability distribution for the flux of every source (Hurley et al., 2017). This approach has several advantages. First, it provides uncertainty quantification by naturally capturing the credible intervals and degeneracies that arise when multiple sources are close together. Second, the joint posterior reveals covariance between sources, quantifying how the flux density estimate of one source is correlated with its neighbours. Third, the Bayesian framework correctly models the behavior of sources at or below the noise level.
The core assumption of XID+ is that the observed map data, , can be modelled as a linear combination of the flux densities, , of a known set of prior sources, each treated as a point source convolved with the instrument’s Point Spread Function (PSF), plus a background () and noise component (). For this work, we use the Gaussian restoring beam from the low-resolution MIGHTEE data. This noise is modelled as a Gaussian that accounts for both thermal noise and confusion noise arising from the blending of faint unresolved sources (e.g.,Condon 1974). For a map flattened into a vector of pixels, this can be expressed as:
@@
Here, is the likelihood of observing the data given a set of fluxes, and is the prior probability distribution of the fluxes. The likelihood is modelled as a multivariate Gaussian distribution based on the noise properties of the map. The standard XID+ framework calculates the total variance by adding the thermal variance (from a user-provided map) and a fitted confusion variance in quadrature (). However, a pure thermal noise map was not available for the MIGHTEE DR1 data products; instead, a total noise RMS map, which already accounts for both effects, was provided (Hale et al., 2025). This means the standard likelihood calculation within XID+ could not be used without modification.
@@
For the deblending performed in this work, we use a simple, uninformative (flat) prior for the flux of each source, typically bounded between 0 and a sufficiently high value, to avoid biasing the resulting measurements towards any pre-conceived model (Hurley et al., 2017; Wang et al., 2024a). The role of the ancillary data is therefore to define an optimal and pure list of source positions, rather than to inform the flux priors directly.
@@
Solving for the full posterior distribution in Equation 2 is a high-dimensional problem that is computationally intractable to solve analytically. XID+ addresses this by using the probabilistic programming language Stan (Carpenter et al., 2017). Specifically, it uses the No-U-Turn Sampler (NUTS, Hoffman and Gelman 2014), to efficiently draw samples from the high-dimensional posterior distribution. To accommodate the MIGHTEE DR1 noise map, we modified the XID+ Stan code to use the pre-computed total variance from the DR1 RMS map at each pixel () directly in the likelihood calculation. While this means we do not derive a separate confusion noise component, it ensures the deblending is performed using the most accurate, position-dependent noise estimate available. An important parameter governing the sampler’s exploration is max_treedepth, which controls the complexity of the integration paths, with a default value of 10. Testing on real MIGHTEE data in Section 5.3 shows that increasing this to 15 was required for convergence. Following the recommendation of Hurley et al. (2017), we run four independent MCMC chains to ensure convergence.
@@
Fitting the whole map at once is computationally unfeasible, so the analysis is partitioned into small tiles using the HEALPix framework (Górski et al., 2005), each 3.4 arcmin across (HEALPix order 10). However, to ensure an accurate flux measurement for sources within one of these small tiles, the deblending fit must account for flux spilling in from bright sources in the surrounding area. Therefore, the fit is performed on a much larger ‘fitting’ tile, 27.5 arcmin across (HEALPix order 7), which fully encompasses the smaller results tile. Once the fit on the larger tile is complete, only the results for sources within the central small tile are retained. The primary output for each source within a successfully fitted tile is a chain of samples from its posterior distribution. The point estimates for the deblended flux densities for each source are taken as the median (50th percentile) of its posterior distribution for each source. The final catalogue is produced by repeating this process for all tiles across the map and combining the results.